OpenStudy (_trilogy):

Celia filled the glasses shown below completely with water. The total amount of water that Celia poured into the glasses is 80 cubic centimeters. What is the height of glass 1? Round your answer to the nearest tenth. (Use π = 3.14.) Note that all measurements are in centimeters and images are not drawn to scale. A cylinder with width 3 and height unknown is labeled glass 1, and a cone with height 6.3 and width 4 is labeled glass 2.

OpenStudy (dumbcow):

Volume of cylinder = \[\pi r^2 * h\] r = 3/2 V = 80 solve for height